OpenStudy (anonymous):

show that sin^2x-in^2y=sin(x+y)sin(x-y)

OpenStudy (kc_kennylau):

try expanding the rhs by \(\sin(\theta+\phi)=\sin\theta\cos\phi+\cos\theta\sin\phi\) ?

OpenStudy (anonymous):

how would i use that

OpenStudy (anonymous):

how come i wouldn't use sin(a-b)sin(a)cos(b)-cos(a)sin(b) instead

OpenStudy (kc_kennylau):

@Kainui sorry I do not know how to solve this problem

OpenStudy (anonymous):

its ok do you know how to solve this one csc-1 (2)

OpenStudy (kc_kennylau):

\[\begin{array}{rl}x&=&\csc^{-1}2\\\csc x&=&2\\\sin x&=&0.5\end{array}\]

OpenStudy (kc_kennylau):

gtg bye :)

OpenStudy (anonymous):

let x=u+v, y=u-v sin(u+v)^2-sin(u-v)^2=sin(2u)sin(2v) (sin(u+v)+sin(u-v))(sin(u+v)-sin(u-v))=sin(2u)sin(2v) (2sin(u)cos(v))(2cos(u)sin(v))=sin(2u)sin(2v) (2sin(u)cos(u))(2sin(v)cos(v))=sin(2u)sin(2v)

OpenStudy (anonymous):

2sin(u)cos(u)=sin(2u) 2sin(v)cos(v)=sin(2v)